# Dream, Lift, Animate: From Single Images to Animatable Gaussian Avatars

### v1

@@ B.1 Supplementary Comparisons
We provide supplementary comparisons with the best performing and concurrent work IDOL [77] on 4D-Dress [60]. Tbl. 3 reports metrics, and Fig. 9 shows visual results for novel view synthesis. We complement these results with a comparison for animation in Fig. 10. Please see the supplementary video for in-the-wild results on SHHQ [27].
@@ B.2 Supplementary Ablations
Table 4: Ablation studies (Sec. B.2) Ours* denotes our model after finetuning on SHHQ.
@@ C. Inputs
We complement the ablations in the main paper with detailed metrics in Tbl. 4 and visuals in Fig. 11. The first set of ablations directly projects to the UV space with the SMPL-X mesh without learning the UV mapping (A.i). The second set replaces the vertex position map with a learned query, which is optimized during training (A.ii). When conditionals are missing in the Gaussian Parameter Decoder, the performance degrades, and we observe artifacts close to the surface (B). Finally, we simulate noisy SMLP-X fits (C.i), purely synthesized inputs (C.ii), and the availability of ground truth views (C.iii). The row Ours* indicates our model finetuned on SHHQ.

### v2

@@ B.1 Supplementary Comparisons
We extend the visual results from the main paper with additional examples from ActorsHQ [28] in Fig. 9 and 10, and we provide more examples from in-the-wild SHHQ images in Fig. 8. Tbl. 3 provides supplementary comparisons for novel pose synthesis on ActorsHQ. In addition, we compare with with the best-performing related work, IDOL [80], on 4D-Dress [63]. Tbl. 4 reports metrics, and Fig. 11 shows visual results for novel view synthesis. We complement these results with a comparison for animation in Fig. 12. Finally, Fig. 15 provides a high-resolution variant of the application figure in the main paper. Please see the supplementary video for in-the-wild results on SHHQ [28].
@@ C. Inputs
Table 5: Supplementary ablation studies (Sec. B.2). Ours* denotes our model after finetuning on SHHQ.
We complement the ablations in the main paper with detailed metrics in Tbl. 5 and visuals in Fig. 16. The first set of ablations directly projects to the UV space with the SMPL-X mesh without learning the UV mapping (A.i). The second set replaces the vertex position map with a learned query, which is optimized during training (A.ii). When conditionals are missing in the Gaussian Parameter Decoder, the performance degrades, and we observe artifacts close to the surface (B). Finally, we simulate noisy SMLP-X fits (C.i), purely synthesized inputs (C.ii), and the availability of ground truth views (C.iii). The row Ours* indicates our model finetuned on SHHQ.
State-of-the art methods [80, 49] ignore pose-dependent effects and model view-dependent effects with spherical harmonics [32]. The design of our Gaussian Parameter Decoder (GPD, Sec. 3.3 and Fig. 4 in the main paper) enables pose- and view-dependent effects by conditioning on a UV-space vertex position map, surface normals, and Plücker rays. Fig. 14 visualizes pose-dependent effects by conditioning the Gaussian Parameter Decoder on vertex position maps and surface normals from other samples, and Fig. 13 shows how the rendering changes when feeding different view directions as conditional inputs to the GPD.